OpenStudy (anonymous):

Will fan and medal please help in 8th grade pre-algebra

OpenStudy (anonymous):

Which of these is a simplified form of the equation 8y + 4 = 6 + 2y + 1y? 5y = 2 5y = 10 11y = 2 11y = 10

HanAkoSolo (jamierox4ev3r):

Just add the like terms :) 8y+4=6+2y+1y Can you identify the like terms for me? Hint: 6 and 4 are like terms

OpenStudy (anonymous):

will it be B or D???

OpenStudy (callielovesyhuu):

a

HanAkoSolo (jamierox4ev3r):

@callielovesyhuu don't provide direct answers. but no, neither B or D is correct

OpenStudy (anonymous):

So I will do 6-4 and that gives me 2 correct???

HanAkoSolo (jamierox4ev3r):

literally, all you have to do is this 8y+4=6+2y+1y 8y+4=6+3y -3y -3y ---------- 5y+4=6 -4 -4 ------- 5y=2 ^ make sense?

OpenStudy (anonymous):

Yes I understand thanks

OpenStudy (anonymous):

Can I ask 1 more question

HanAkoSolo (jamierox4ev3r):

sure :)

OpenStudy (anonymous):

What is the value of p in the equation 4p + 3(p + 4) = 12p + 12? 12 1 0 −1

HanAkoSolo (jamierox4ev3r):

can you please open another question? this thread is getting a little laggy for me lol

OpenStudy (anonymous):

sure